# Cohomology and Deformations of left-symmetric Rinehart Algebras

### v3

@@ Abstract
We introduce a notion of left-symmetric Rinehart algebras, which is a generalization of a left-symmetric algebras. The left multiplication gives rise to a representation of the corresponding sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebra from $\mathcal{O}$-operators on Lie-Rinehart algebra. We extensively investigate representations of a left-symmetric Rinehart algebras. Moreover, We construct a graded Lie algebra on the space of multi-derivations whose Maurer–Cartan elements characterize left-symmetric Rinehart algebras and study deformations of left-symmetric Rinehart algebras, which is controlled by the second cohomology class in the deformation cohomology. We also give the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebra.
Keywords: left-symmetric Rinehart algebras, representations, graded Lie algebras, Maurer-Cartan element, cohomology, deformations, Nijenhuis operators.
MSC(2020): Primary 17E05, 53D17, 17B70; Secondary 14B12, 06B15
@@ 1 Introduction
Left-symmetric algebras are algebras for which the associator $(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z)$ satisfies the identity $(x,y,z)=(y,x,z)$. These algebras appeared as early as 1896 in the work of Cayley [7] as rooted tree algebras. In the 1960s, they also arose from the study of several topics in geometry and algebra, such as convex homogenous cones [32], affine manifolds and affine structures on Lie groups [20, 29] and deformation of associative algebras [19]. In 2006, Burde [5] wrote an interesting survey showing the importance of left-symmetric algebras in many areas, such as vector fields, rooted tree algebras, vertex algebras, operad theory, deformation complexes of algebras, convex homogeneous cones, affine manifolds and left-invariant affine structures on Lie groups [5].
Left symmetric algebras are the underlying algebraic structures of non-abelian phase spaces of Lie algebras [21, 1], leading to a bialgebra theory of left-symmetric algebras [3]. They can also be seen as the algebraic structures behind the classical Yang-Baxter equations. Precisely they provide a construction of solutions of the classical Yang-Baxter equations in certain semidirect product Lie algebra structures (that is, over the double spaces) induced by left-symmetric algebras [22, 2].
The notion of Lie-Rinehart algebras was introduced by J. Herz in [14] and further developed in [30, 31]. A a notion of (Poincaré) duality for this class of algebras was introduced in [17, 18]. Lie-Rinehart structures have been the subject of extensive studies, in relations to symplectic geometry, Poisson structures, Lie groupoids and algebroids and other kind of quantizations (see [14, 16, 23, 24, 25, 26, 27]). For further details and a history of the notion of Lie-Rinehart algebra, we refer the reader to [16]. Lie-Rinehart algebras have been investigated furthermore in [4, 8, 10].
@@
In this paper, we introduce a notion of a left-symmetric Rinehart algebra, which is a generalization of a left-symmetric algebra and an algebraic version of left symmetric algebroid. The following diagram shows how left-symmetric Rinehart algebras fit in relation to Lie algebras, left-symmetric algebras and Lie-Rinehart algebras.
Lie algebra commutatorgeneralizationLie-Rinehart algebracommutatorLeft-symmetric algebraRota-Baxter operatorgeneralizationLeft-symmetric Rinehart Rota-Baxter operator
The paper is organized as follows. In Section 2, we recall some definitions concerning left-symmetric algebras and Lie-Rinehart algebra. In Section 3, we introduce the notion of left-symmetric Rinehart algebra and give some of its properties. As in the case of a left-symmetric algebra, one can obtain the sub-adjacent Lie-Rinehart from a left-symmetric Rinehart by using the commutator. The left multiplication gives rise to a representation of the sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebra using $\mathcal{O}$-operators. Section 4 is devoted to the study of representations and cohomology of left-symmetric Rinehart algebras. In Section 5, we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which gives rises a coboundary operator. Section 6 is devoted to introduce the deformation cohomology associated to a left-symmetric Rinehart algebra, which controls the deformations. In Section 7, we introduce the notion of a Nijenhuis operator, which could generate a trivial deformation. In addition, we investigate some connection between $\mathcal{O}$-operator and Nijenhuis operator.
@@ 2 Preliminaries
$\rho(x)\circ\mu(y)-\mu(y)\circ\rho(x)=\mu(x\cdot y)-\mu(y)\circ\mu(x),\quad\forall~{}x,y\in L.$ (2.1)
@@
$C^{n+1}(L,M)=\mathrm{Hom}(\wedge^{n}L\otimes L,M),\quad\forall n\geq 0.$ (2.2)
@@
$\delta\omega(x_{1},x_{2},\dots,x_{n+1})$
$=$ $\sum_{i=1}^{n}(-1)^{i+1}\rho(x_{i})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n+1})$
$+\sum_{i=1}^{n}(-1)^{i+1}\mu(x_{n+1})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i})$
$-\sum_{i=1}^{n}(-1)^{i+1}\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i}\cdot x_{n+1})$
$+\sum_{1\leq i<j\leq n}(-1)^{i+j}\omega([x_{i},x_{j}],x_{1},\dots,\widehat{x_{i}},\dots,\widehat{x_{j}},\dots,x_{n+1}).$
We then have the following lemma whose proof comes from a direct computation using identity (2.1).
Lemma 2.4.
[6] The map $\delta$ satisfies $\delta^{2}=0$.
@@
    $\rho$ is a representation of $(L,[\cdot,\cdot])$ on $A$.
@@
    $[x,ay]=\rho(x)ay+a[x,y],~{}~{}~{}\forall a\in A,x,y\in L.$ (2.3)
Let $(L,A,[\cdot,\cdot]_{L},\rho)$ and $(L^{\prime},A^{\prime},[\cdot,\cdot]_{L^{\prime}},\rho^{\prime})$ be two Lie-Rinehart algebras, then a Lie-Rinehart algebra homomorphism is defined as a pair of maps $(g,f)$, where the maps $f:L\rightarrow L^{\prime}$ and $g:A\rightarrow A^{\prime}$ are two $\mathbb{K}$-algebra homomorphisms such that:
@@
Let $M$ be an $A$-module. Then $M$ is a module over a Lie-Rinehart algebra $(L,A,[\cdot,\cdot],\rho)$ if there exits a map $\theta:L\otimes M\rightarrow M$ such that:
@@
We have the following lemma giving a characterization of when $\theta$ is a representation.
@@
$\theta$ is representation if and only if $L\oplus M$ is Lie-Rinehart algebra over $A$.
3 Some basic properties of a left-symmetric Rinehart algebra
In this section, we introduce a notion of left-symmetric Rinehart algebras illustrated by some examples. As in the case of a left-symmetric algebra, we obtain the sub-adjacent Lie-Rinehart from a left-symmetric Rinehart using the commutator. In addition, we construct left-symmetric Rinehart algebra using $\mathcal{O}$-operators.
@@ 3 Some basic properties of a left-symmetric Rinehart algebra
    $\ell$ is a left representation of $(L,\cdot)$ on $A$.
@@
    $x\cdot(ay)=$ $\ell(x)ay+a(x\cdot y),$ (3.1)
    $(ax)\cdot y=$ $a(x\cdot y).$ (3.2)
Remark 3.2.
If $\ell$ is surjective (respectively bijective) then the left-symmetric Rinehart algebra $(L,A,\cdot,\ell)$ is called transitive (respectively regular) left-symmetric Rinehart algebra.
@@
A Novikov Poisson algebra is a left-symmetric Rinehart algebra see [35]).
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and let $L\oplus A=\{x+a~{}|~{}x\in L,a\in A\}$ be the direct sum of $L$ and $A$. Then $(L\oplus A,A,\cdot_{L\oplus A},\ell_{L\oplus A})$ is a left-symmetric Rinehart algebra, where the $\cdot_{L\oplus A}$ is defined by, for all $x_{1},x_{2}\in L$, $a_{1},a_{2}\in A$;
$\begin{split}(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})=&x_{1}\cdot x_{2}+\ell(x_{1})(a_{2});\end{split}$
and $\ell_{L\oplus A}:L\oplus A\rightarrow Der(A)$ is defined by $\ell_{L\oplus A}(a_{1}+x_{1})=\ell(x_{1})$. Indeed, it obvious to show that $(L\oplus A,\cdot_{L\oplus A})$ is a left-symmetric algebra, $\ell_{L\oplus A}$ is a representation of left-symmetric algebra $L\oplus A$ and $\ell_{L\oplus A}\in Der(A)$.
By direct calculation, we have $\ell_{L\oplus A}(b(x_{1}+a_{1}))=b\ell_{L\oplus A}(x_{1}+a_{1})$ for all $b,a_{1}\in A$ and $x_{1}\in L$. On the other hand, let $x_{1},x_{2}\in L$ and $b,a_{1},a_{2}\in A$ we have
@@
Definition 3.6.
A subalgebra $(S,A)$ of $(L,A,\cdot,\ell)$, $S$ for short, is a left-symmetric subalgebra of $L$ such that $AS\subset S$ and $S$ acts on $A$ via the composition $S\hookrightarrow L\rightarrow Der(A)$.
A subalgebra $(I,A)$, $I$ for short, of $L$ is called an ideal if $I$ is a left-symmetric ideal of $L$ such that $\ell(I)AL\subset I$.
Now we have the following theorem
@@
Proof. Since $(L,\cdot)$ is a left-symmetric algebra, we have that $(L,[\cdot,\cdot])$ is a Lie algebra. For any $a\in A$, by direct computations, we have
@@
Consider the semidirect product Lie-Rinehart algebra $(L\ltimes_{\theta}M,A,[\cdot,\cdot]_{L\ltimes_{\theta}M},\rho_{L\ltimes_{\theta}M})$, where $\rho_{L\ltimes_{\theta}M}(x+u):=\rho(x)(u)$ and the bracket $[\cdot,\cdot]_{L\ltimes_{\theta}M}$ is given by
@@
$[\tilde{T}(x+u),\tilde{T}(y+v)]_{L\ltimes_{\theta}M}=\tilde{T}([\tilde{T}(x+u),y+v]_{L\ltimes_{\theta}M}+[x+u,\tilde{T}(y+v)]_{L\ltimes_{\theta}M}-\tilde{T}[x+u,y+v]_{L\ltimes_{\theta}M}).$
@@
$u\cdot_{T}v=\theta(T(u))(v),\forall u,v\in M.$
We then have the following proposition
@@
Proof. It is easy to see that $(M,\cdot_{T})$ is a left-symmetric algebra. For any $a\in A$, using Definition 3.1 and Eq.(3.3) we have
@@ 4 Representations of left-symmetric Rinehart algebras
In this section, we develop the notion of representations of a left-symmetric Rinehart algebra and give a cohomology theory with the coefficient in a representation.
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $M$ be an $A$-module. A representation of $A$ on $M$ consists of a pair $(\rho,\mu)$, where $\rho:L\otimes M\rightarrow M$ is a left representation of $(L,\cdot)$ on $M$ and $\mu:M\otimes L\rightarrow M$ is a linear map, such that for all $x,y\in L$ and $\ m\in M$, we have
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra. It is obvious that if $(M,\rho)$ is a representation of the sub-adjacent Lie-Rinehart algebra $(L,A,[\cdot,\cdot]^{C},\ell)$, then $(M;\rho,0)$ is a representation of the left-symmetric Rinehart algebras $(L,A,\cdot,\ell)$.
@@
$(x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})$ $=$ $x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1},$ (4.2)
$\ell_{L\oplus M}(x_{1}+m_{1})$ $=$ $\ell(x_{1}),$ (4.3)
@@
Proof. Let $(M;\rho,\mu)$ be a representation. It is straightforward to see that $(L\oplus M,A,\cdot_{L\oplus M})$ is a left-symmetric algebra. For any $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$, we have
@@
$\langle\rho^{*}(x)\xi,m\rangle=\ell(x)\langle m,\xi\rangle-\langle\rho(x)m,\xi\rangle\ \ \textrm{and}\ \ \langle\mu^{*}(x)\xi,m\rangle=-\langle\mu(x)m,\xi\rangle,$
@@
Proof. Since $(M;\rho,\mu)$ is a representation of the left-symmetric algebra $(L,A,\cdot,\ell)$, using Proposition 4.2 we have that $(L\oplus M,A,\cdot_{L\oplus M},\ell_{L\oplus M})$ is a left-symmetric Rinehart algebra. Consider its sub-adjacent Lie-Rinehart algebra $(L\oplus M,A,\cdot_{L\oplus M},[\cdot,\cdot]_{L\oplus M},\ell_{L\oplus M})$, we have
@@
$-\langle(\rho^{*}-\mu^{*})(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(y)(\rho^{*}-\mu^{*})(x)\xi,m\rangle$ $=$ $\langle\mu^{*}(y)\xi,(\rho-\mu)(x)m\rangle-\langle(\rho^{*}-\mu^{*})(x)\xi,\mu(y)m\rangle$
$=$ $-\langle\xi,\mu(y)(\rho-\mu)(x)m\rangle+\langle\xi,(\rho-\mu)(x)\mu(y)m\rangle$
$=$ $\langle\xi,(\rho-\mu)(x)\mu(y)m-\mu(y)(\rho-\mu)(x)m\rangle$
$=$ $\langle\xi,\mu(x\cdot y)m-\mu(x)\mu(y)m\rangle$
$=$ $\langle-\mu^{*}(x\cdot y)\xi-\mu^{*}(y)\mu^{*}(x)\xi,m\rangle.$
@@
    The left-symmetric Rinehart algebra $L\ltimes_{\rho,\mu}M$ and $L\ltimes_{\rho-\mu,0}M$ have the same sub-adjacent Lie-Rinehart algebra $L\ltimes_{\rho-\mu}M.$
@@
Proof. $(1)\Rightarrow(2)$ Using Proposition 4.3, since $(M;\rho,\mu)$ is representation of $L$ then $(M^{*},\rho*,\mu^{*})$ is representation of $L$.
$(2)\Rightarrow(3)$ For any $x,y\in L$, $\xi\in M^{*}$ and $m\in M$ we have
$\langle\xi,\mu(x)\mu(y)m\rangle-\langle\xi,\mu(y)\mu(x)m\rangle$ $=$ $\langle\xi,\mu(x)\mu(y)m\rangle-\langle\xi,\mu(x\cdot y)m\rangle+\langle\xi,\rho(x)\mu(y)m\rangle$
$-\langle\xi,\mu(y)\rho(x)m\rangle$
$=$ $\langle\mu^{*}(y)\mu^{*}(x)\xi,m\rangle+\langle\mu^{*}(x\cdot y)\xi,m\rangle+\langle\mu^{*}(y)\rho^{*}(x)\xi,m\rangle$
$-\rho^{*}(x)\langle\mu^{*}(y)\xi,m\rangle$
$=$ $0$
Using Eq. (4.1) it is easy to show that $(3)\Rightarrow(1)$.
Example 4.6.
Let $(L,A,\cdot,\ell)$ be a transitive left-symmetric Rinehart algebra and let $K=\ker(\ell)$. Then we have
  • $\rm(1)$
    $(L,\mathrm{ad}^{L},0)$ is a representation and $(L^{*},ad^{L^{*}},0)$ is a representation.
  • $\rm(2)$
    $(K,\mathrm{ad}^{L}-\mathrm{ad}^{R},0)$ is a representation of $L$ on $K=\ker(\ell)$ and $(K^{*},{\mathrm{ad}}^{L^{*}},0)$ is a representation of $L$ on $K^{*}$.
  • $\rm(3)$
    If $K$ is an ideal of $L$, then $(K,\mathrm{ad}^{L},\mathrm{ad}^{R})$ is a representation of $L$ on $K$ and $(K^{*},\mathrm{ad}^{L^{*}}-\mathrm{ad}^{R^{*}},-\mathrm{ad}^{R^{*}})$ is also a representation on $K^{*}$ by Proposition 4.3.

### v5

@@ Abstract
We introduce a notion of left-symmetric Rinehart algebras, which is a generalization of the notion of left-symmetric algebras. The left multiplication gives rise to a representation of the corresponding sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebras from $\mathcal{O}$-operators on Lie-Rinehart algebras. We extensively investigate representations of left-symmetric Rinehart algebras. Moreover, we construct a graded Lie algebra on the space of multi-derivations whose Maurer–Cartan elements characterize left-symmetric Rinehart algebras and study deformations of left-symmetric Rinehart algebras, which are controlled by the second cohomology class in the deformation cohomology. We also give the relationships between $\mathcal{O}$-operators and Nijenhuis operators on left-symmetric Rinehart algebras.
keywords:
left-symmetric Rinehart algebra, representation, graded Lie algebra, Maurer-Cartan element, cohomology, deformation, Nijenhuis operator.
[…]
@@ 1 Introduction
Left-symmetric algebras are algebras for which the associator
$(x,y,z):=(x\cdot y)\cdot z-x\cdot(y\cdot z)$
satisfies the identity $(x,y,z)=(y,x,z)$. These algebras appeared as early as 1896 in the work of Cayley [8] as rooted tree algebras. In the 1960s, they also arose from the study of several topics in geometry and algebra, such as convex homogenous cones [30], affine manifolds and affine structures on Lie groups [19, 27] and deformations of associative algebras [14]. In 2006, Burde [6] wrote an interesting survey showing the importance of left-symmetric algebras in many areas, such as vector fields, rooted tree algebras, vertex algebras, operad theory, deformation complexes of algebras, convex homogeneous cones, affine manifolds and left-invariant affine structures on Lie groups [6].
Left symmetric algebras are the underlying algebraic structures of non-abelian phase spaces of Lie algebras [21, 1], leading to a bialgebra theory of left-symmetric algebras [3]. They can also be seen as the algebraic structures behind the classical Yang-Baxter equations. Precisely, they provide a construction of solutions of the classical Yang-Baxter equations in certain semidirect product Lie algebra structures (that is, over the double spaces) induced by left-symmetric algebras [22, 2].
The notion of Lie-Rinehart algebras was introduced by J. Herz in [15] and further developed in [28, 29]. A a notion of (Poincaré) duality for this class of algebras was introduced in [16, 17]. Lie-Rinehart structures have been the subject of extensive studies, in relations to symplectic geometry, Poisson structures, Lie groupoids and algebroids and other kinds of quantizations (see [18, 20, 25, 24, 23, 26]). For further details and a history of the notion of Lie-Rinehart algebra, we refer the reader to [18]. Lie-Rinehart algebras have been investigated furthermore in [4, 7, 12, 11].
@@
In this paper, we introduce a notion of left-symmetric Rinehart algebras, which is a generalization of a left-symmetric algebra and an algebraic version of left symmetric algebroids. The following diagram shows how left-symmetric Rinehart algebras fit in relation to Lie algebras, left-symmetric algebras and Lie-Rinehart algebras.
Lie algebrageneralizationLie-Rinehart algebraLeft-symmetric algebrageneralizationLeft-symmetric Rinehart
The paper is organized as follows. In Section 2, we recall some definitions concerning left-symmetric algebras and Lie-Rinehart algebra. In Section 3, we introduce the notion of left-symmetric Rinehart algebra and give some of its properties. As in the case of a left-symmetric algebras, one can obtain the sub-adjacent Lie-Rinehart algebra from a left-symmetric Rinehart algebra by using the commutator. The left multiplication gives rise to a representation of the sub-adjacent Lie-Rinehart algebra. We construct left-symmetric Rinehart algebras using $\mathcal{O}$-operators. Section 4 is devoted to the study of representations and cohomology of left-symmetric Rinehart algebra. In Section 5, we construct a graded Lie algebra whose Maurer-Cartan elements are left-symmetric Rinehart algebras which give rise to a coboundary operator. Section 6 is devoted to introduce the deformation cohomology associated to a left-symmetric Rinehart algebra, which controls the deformations. In Section 7, we introduce the notion of a Nijenhuis operator, which could generate a trivial deformation. In addition, we investigate some connection between $\mathcal{O}$-operators and Nijenhuis operators.
@@ 2 Preliminaries
$\rho(x)\circ\mu(y)-\mu(y)\circ\rho(x)=\mu(x\cdot y)-\mu(y)\circ\mu(x),\quad\forall~{}x,y\in L.$ (1)
@@
$C^{n+1}(L,M)=\mathrm{Hom}(\wedge^{n}L\otimes L,M),\quad\forall n\geq 0.$ (2)
@@
$\delta\omega(x_{1},x_{2},\dots,x_{n+1})$ $=$ $\sum_{i=1}^{n}(-1)^{i+1}\rho(x_{i})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n+1})$
$\quad+\sum_{i=1}^{n}(-1)^{i+1}\mu(x_{n+1})\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i})$
$\quad-\sum_{i=1}^{n}(-1)^{i+1}\omega(x_{1},\dots,\widehat{x_{i}},\dots,x_{n},x_{i}\cdot x_{n+1})$
$\quad+\sum_{1\leq i<j\leq n}(-1)^{i+j}\omega([x_{i},x_{j}],x_{1},\dots,\widehat{x_{i}},\dots,\widehat{x_{j}},\dots,x_{n+1}).$
We then have the following lemma whose proof comes from a direct computation using identity (1).
Lemma 2.4 (See [5]).
The map $\delta$ satisfies $\delta^{2}=0$.
@@
    For all $a\in A$ and $x,y\in L$
    $\rho([x,y])=\rho(x)\rho(y)-\rho(x)\rho(y)\quad\text{and}\quad\rho(ax)=a\rho(x).$
@@
    $[x,ay]=\rho(x)ay+a[x,y],~{}~{}~{}\forall a\in A,x,y\in L.$ (3)
Let $(L,A,[\cdot,\cdot]_{L},\rho)$ and $(L^{\prime},A^{\prime},[\cdot,\cdot]_{L^{\prime}},\rho^{\prime})$ be two Lie-Rinehart algebras, then a Lie-Rinehart algebra homomorphism is defined as a pair of maps $(g,f)$, where the maps $f:L\rightarrow L^{\prime}$ and $g:A\rightarrow A^{\prime}$ are two algebra homomorphisms such that:
@@
Let $M$ be an $A$-module. Then $M$ is a module over a Lie-Rinehart algebra $(L,A,[\cdot,\cdot],\rho)$ if there exists a map $\theta:L\otimes M\rightarrow M$ such that:
@@
We have the following lemma giving a characterization of of the $\theta$ which are representations.
@@
The map $\theta$ is representation if and only if $L\oplus M$ is Lie-Rinehart algebra over $A$, where $[\cdot,\cdot]_{L\oplus M}$ and $\theta_{L\oplus M}$ are given by
$[x_{1}+m_{1},x_{2}+m_{2}]_{L\oplus M}$ $=$ $[x_{1},x_{2}]+\rho(x_{1})m_{2}-\rho(x_{2})m_{1},$
$\theta_{L\oplus M}(x_{1}+m_{1})$ $=$ $\theta(x_{1})$
for all $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$.
3 Some basic properties of a left-symmetric Rinehart algebras
In this section, we introduce a notion of left-symmetric Rinehart algebras illustrated by some examples. As in the case of a left-symmetric algebra, we obtain the sub-adjacent Lie-Rinehart algebra from a left-symmetric Rinehart algebra using the commutator. In addition, we construct left-symmetric Rinehart algebras using $\mathcal{O}$-operators.
@@ 3 Some basic properties of a left-symmetric Rinehart algebras
    For all $a\in A$ and $x,y\in L$
    $\ell(x\cdot y-y\cdot x)=\ell(x)\ell(y)-\ell(x)\ell(y)\quad\text{and}\quad\ell(ax)=a\ell(x).$
@@
    $x\cdot(ay)=$ $\ell(x)ay+a(x\cdot y),$ (4)
    $(ax)\cdot y=$ $a(x\cdot y).$ (5)
@@
A Novikov Poisson algebra is a left-symmetric Rinehart algebra (see [32]).
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and let $L\oplus A$be the direct sum of $L$ and $A$. Then $(L\oplus A,A,\cdot_{L\oplus A},\ell_{L\oplus A})$ is a left-symmetric Rinehart algebra, where the $\cdot_{L\oplus A}$ is defined by the following expression, for all $x_{1},x_{2}\in L$, $a_{1},a_{2}\in A$;
$(x_{1}+a_{1})\cdot_{L\oplus A}(x_{2}+a_{2})=x_{1}\cdot x_{2}+\ell(x_{1})(a_{2});$
and $\ell_{L\oplus A}:L\oplus A\rightarrow Der(A)$ is defined by $\ell_{L\oplus A}(a_{1}+x_{1})=\ell(x_{1})$. Indeed, it obvious that $(L\oplus A,\cdot_{L\oplus A})$ is a left-symmetric algebra, $\ell_{L\oplus A}$ is a representation of left-symmetric algebra $L\oplus A$ and $\ell_{L\oplus A}\in Der(A)$.
By direct calculation, we have $\ell_{L\oplus A}(b(x_{1}+a_{1}))=b\ell_{L\oplus A}(x_{1}+a_{1})$ for all $b,a_{1}\in A$ and $x_{1}\in L$. On the other hand, letting $x_{1},x_{2}\in L$ and $b,a_{1},a_{2}\in A$, we have
@@
Now we have the following theorem.
@@
Proof 3.6.
Since $(L,\cdot)$ is a left-symmetric algebra, we have that $(L,[\cdot,\cdot])$ is a Lie algebra. For any $a\in A$, by direct computations, we have
@@
Consider the semidirect product Lie-Rinehart algebra
$(L\ltimes_{\theta}M,A,[\cdot,\cdot]_{L\ltimes_{\theta}M},\rho_{L\ltimes_{\theta}M}),$
where $\rho_{L\ltimes_{\theta}M}(x+u):=\rho(x)(u)$ and the bracket $[\cdot,\cdot]_{L\ltimes_{\theta}M}$ is given by
@@
$[\tilde{T}(x+u),\tilde{T}(y+v)]_{L\ltimes_{\theta}M}=\tilde{T}\Big{(}[\tilde{T}(x+u),y+v]_{L\ltimes_{\theta}M}+[x+u,\tilde{T}(y+v)]_{L\ltimes_{\theta}M}-\tilde{T}[x+u,y+v]_{L\ltimes_{\theta}M}\Big{)}.$
@@
$u\cdot_{T}v=\theta(T(u))(v),\>\forall u,v\in M.$
We then have the following proposition.
@@
Proof 3.12.
It is easy to see that $(M,\cdot_{T})$ is a left-symmetric algebra. For any $a\in A$, using Definition 3.1 and equation (6) we have
@@ 4 Representations of left-symmetric Rinehart algebras
In this section, we develop the notion of representations of a left-symmetric Rinehart algebra and give a cohomology theory with coefficients in a representation.
@@
Let $(L,A,\cdot,\ell)$ be a left-symmetric Rinehart algebra and $M$ be an $A$-module. A representation of $A$ on $M$ consists of a pair $(\rho,\mu)$, where $\rho$ is a representation of the sub-adjacent Lie-Rinehart algebra $(L,A,[\cdot,\cdot]^{C},\ell)$ and $\mu:L\rightarrow End(M)$ is a linear map, such that for all $x,y\in L$ and $\ m\in M$, we have
@@
$(x_{1}+m_{1})\cdot_{L\oplus M}(x_{2}+m_{2})$ $=$ $x_{1}\cdot x_{2}+\rho(x_{1})m_{2}+\mu(x_{2})m_{1},$ (9)
$\ell_{L\oplus M}(x_{1}+m_{1})$ $=$ $\ell(x_{1}),$ (10)
@@
Proof 4.3.
Let $(M;\rho,\mu)$ be a representation. It is straightforward to see that $(L\oplus M,A,\cdot_{L\oplus M})$ is a left-symmetric algebra. For any $x_{1},x_{2}\in L$ and $m_{1},m_{2}\in M$, we have
@@
$\langle\rho^{*}(x)\xi,m\rangle=\ell(x)\langle\xi,m\rangle-\langle\rho(x)m,\xi\rangle\ \ \textrm{and}\ \ \langle\mu^{*}(x)\xi,m\rangle=-\langle\xi,\mu(x)m\rangle,$
@@
Proof 4.5.
Since $(M;\rho,\mu)$ is a representation of the left-symmetric algebra $(L,A,\cdot,\ell)$, using Proposition 4.2 we have that $(L\oplus M,A,\cdot_{L\oplus M},\ell_{L\oplus M})$ is a left-symmetric Rinehart algebra. Consider its sub-adjacent Lie-Rinehart algebra $(L\oplus M,A,\cdot_{L\oplus M},[\cdot,\cdot]_{L\oplus M},\ell_{L\oplus M})$. We have
@@
$-\langle(\rho^{*}-\mu^{*})(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(y)(\rho^{*}-\mu^{*})(x)\xi,m\rangle$
$=-\langle\rho^{*}(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(x)\mu^{*}(y)\xi,m\rangle+\langle\mu^{*}(y)\rho^{*}(x)\xi,m\rangle-\langle\mu^{*}(y)\mu^{*}(x)\xi,m\rangle$
$=\ell(x)\langle\xi,\mu(y)m\rangle-\langle\xi,\mu(y)\rho(x)m\rangle+\langle\xi,\mu(y)\mu(x)m\rangle$
$\qquad-\ell(x)\langle\xi,\mu(y)m\rangle+\langle\xi,\rho(x)\mu(y)m\rangle-\langle\xi,\mu(x)\mu(y)m\rangle$
$=\langle\xi,\mu(x\cdot y)m\rangle-\langle\xi,\mu(y)\mu(x)m\rangle+\langle\xi,\mu(y)\mu(x)m\rangle-\langle\xi,\mu(x)\mu(y)m\rangle$
$=\langle(-\mu^{*}(x\cdot y)-\mu^{*}(y)\mu^{*}(x))\xi,m\rangle.$
@@
    The left-symmetric Rinehart algebras $L\ltimes_{\rho,\mu}M$ and $L\ltimes_{\rho-\mu,0}M$ have the same sub-adjacent Lie-Rinehart algebra $L\ltimes_{\rho-\mu}M.$